Evaluate the following integral by converting it to cylindrical coordinates: ∫(-3 to 3)∫(0 to sqrt(9-x^2))∫(0 to 9-x^2-y^2) sqrt(x^2+y^2) dz dy dx


1
Expert's answer
2020-11-17T12:52:43-0500

"I = \int_{-3}^{3}\int_0^{\sqrt{9 - x^2}}\int_0^{\sqrt{9 - x^2 - y^2}} \sqrt{x^2 + y^2}\, \mathrm{d}z\, \mathrm{d}y\, \mathrm{d}x\\

\textsf{Converting to cylindrical coordinates.}\\
\, \mathrm{d}z\, \mathrm{d}y\, \mathrm{d}x = \rho\mathrm{d}z\, \mathrm{d}\rho\, \mathrm{d}\theta\\

x = \rho\cos{\theta}, y = \rho\sin{\theta}\\

x^2 + y^2 = \rho^2(\cos^2{\theta} + \sin^2{\theta}) = \rho^2\\

y = 0, \sin{\theta} = 0, \theta = 2\pi\\

y = \sqrt{9 - x^2}\\

y = \sqrt{9 - \rho^2\cos^2{\theta}}\\

x= -3, \rho\cos{\theta} = -3\\

x= -3, \rho\cos(2\pi) = \rho = -3\\

x= 3, \rho\cos(2\pi) = \rho = 3\\

\therefore y = 3\sin{\theta} = \sqrt{9 - 9\cos^2{2\pi}} = \sqrt{9 - 9} = 0, \theta = 0\\

z = \sqrt{9 - x^2 - y^2} = \sqrt{9 - \rho^2}\\

\begin{aligned}
I = \int_{-3}^{3}\int_0^{\sqrt{9 - x^2}}\int_0^{\sqrt{9 - x^2 - y^2}} \sqrt{x^2 + y^2}\, \mathrm{d}z\, \mathrm{d}y\, \mathrm{d}x &= \int_{-3}^{3}\int_0^{2\pi} \int_0^{\sqrt{9 - \rho^2}} \sqrt{\rho^2} \cdot \rho\mathrm{d}z\, \mathrm{d}\theta\, \mathrm{d}\rho
\\&=\int_{-3}^{3}\int_0^{2\pi} \int_0^{\sqrt{9 - \rho^2}} \rho^2\,\mathrm{d}z\, \mathrm{d}\theta\, \mathrm{d}\rho
\\&=\int_{-3}^{3}\rho^2\,\mathrm{d}\rho\int_0^{\sqrt{9 - \rho^2}} \mathrm{d}z \int_0^{2\pi}\, \mathrm{d}\theta
\\&= 2\pi\int_{-3}^{3}\rho^2\,\mathrm{d}\rho \cdot1 \vert_0^{\sqrt{9 - \rho^2}} 
\\&= 2\pi\int_{-3}^{3}\rho^2\sqrt{9 - \rho^2}\,\mathrm{d}\rho
\end{aligned}\\

\textsf{Substitute}\, \rho = 3\sin\alpha\\
\begin{aligned}
\therefore I &= 2\pi\int_{-\frac{\pi}{2}}^{\frac{\pi}{2}}9\sin^2{\alpha}\sqrt{9 - 9\sin^2\alpha}\cdot3\cos{\alpha}\mathrm{d}\alpha
\\&= 2\pi\int_{0}^{\frac{\pi}{2}}9\sin^2{\alpha}\sqrt{9 - 9\sin^2\alpha}\cdot3\cos{\alpha}\mathrm{d}\alpha
\\&= 324\pi\int_{0}^{\frac{\pi}{2}}\sin^2{\alpha}\cos^2{\alpha}\,\mathrm{d}\alpha
\\&= 324\pi\left(\frac{\alpha}{8} + \frac{\sin{\alpha}\cos{\alpha}}{8}- \frac{\sin{\alpha}\cos^{3}{\alpha}}{4}\right)\vert_{0}^{\frac{\pi}{2}}
\\&= 324\left(\frac{\pi}{16}\right) = \frac{81\pi^2}{4}
\end{aligned}\\


\textsf{Note:}\, \int_{-a}^{a} f(x)\, \mathrm{d}x = 2\int_{0}^{a} f(x)\, \mathrm{d}x, \,\,\textsf{if }\,f(x) \textsf{is even.}"